OpenStudy (anonymous):

Evaluate: tan105deg using the sum and difference formulas. I

OpenStudy (anonymous):

tan(a+b)=(tan a +tan b)/(1-tan a tan b)....put a =60 and b =45 and solve

OpenStudy (anonymous):

where u got 60 and 45?

OpenStudy (anonymous):

tan(60+45)=tan105...right

OpenStudy (anonymous):

can u do it nw

OpenStudy (anonymous):

tan 60 =\[\sqrt{3}\] and tan 45 =1

OpenStudy (anonymous):

i am getting stuck at the part where it is: